OpenStudy (anonymous):

order of operations to work the problem 3-5{-5(7-7)-2(5-3)}

OpenStudy (anonymous):

i worked it out got 10

OpenStudy (anonymous):

Parenthesis first followed by {} then solve. \[3-5{-5(7-7)-2(5-3)} \] so -2(5-3)=-4 so then -5(7-7)=0 then -5{-4}=20 then 3+20=23 So your answer is 23.

OpenStudy (anonymous):

i reworked it and made sure i did symbols first thats were i messed up. thank you i didt work from left to right. i got 23 when i did it right

OpenStudy (anonymous):

Yeah always start on the inside. I did the ( ) to the right first and then did the one to the left which cancelled out the ( ) to the left which narrowed it down. I think all you were forgetting to do was to distribute to get rid of the {( )}.

OpenStudy (anonymous):

yes i didnt multiply the 5 from the (7-7) and the 2 (5-3)